# How to Train a Backdoor-Robust Model on a Poisoned Dataset without Auxiliary Data?

### v1

@@ VI-B2 The Resistance to Adaptive Attacks
To further reveal the potential risk of the AdvrBD, we also consider the adaptive adversaries that try to design special backdoor attacks to escape our AdvrBD method. The adaptive adversaries deliberately establish dependencies between the backdoor and normal functionality, which immensely increases the difficulty of backdoor sample detection. Therefore, we employ three adaptive attacks (TaCT, Adap-Blend, and Adap-Patch) to demonstrate the effectiveness of our defense method. The three adaptive attacks build the dependencies between the backdoor and normal functions by using different poisoning strategies. Table. III shows the defense results of our method and other $7$ state-of-the-art backdoor defense techniques against the $3$ adaptive attacks on $4$ different datasets. From Table III, we have the following three observations. First, our AdvrBD has the lowest ASR (nearly $0.00\%$) while maintaining a satisfactory ACC of the main task. Then, some existing defense methods (CLP and SCAnFT) have a good defense performance against the three adaptive attacks, but they have an obvious decrease in ACC. Specifically, for CLP on the CIFAR10, the ASR is $3.55\%$, and the ACC drops from $83.63\%$ to $52.56\%$. Finally, our method has a better generalization on different datasets. For different datasets, our method always maintains a fine performance, while the performance of the other backdoor defense techniques has an obvious fluctuation. In summary, compared with the $7$ state-of-the-art backdoor defense techniques, our method has a better performance and greater generalization against adaptive attacks.
VI-C Ablation Studies
@@
Different from the existing backdoor defense methods, we also propose to improve the performance of our AdvrBD further by relabeling and relearning the backdoor samples. Hence, in this section, we first show the AdvrBD’s performance after employing relabeling and relearning mechanisms. Then, to demonstrate our AdvrBD’s generalization, we also experimentally explore the performance of our AdvrBD on different partition rates, poison ratios, and models. Finally, we also explore the impact of our AdvrBD on the clean and poisoned dataset.
VI-C1 Performance with Relabeling and Relearning
@@
In our AdvrBD, after identifying the backdoor samples and clean samples, we relabel the backdoor samples and merge them with clean samples to obtain a clean and complete dataset, which is used to fine-tune the clean model then. Accordingly, we conduct experiments on the ResNet18 and CIFAR10 datasets to show the effect of the relabeling and relearning mechanism. Table IV shows the ACC and ASR of our AdvrBD after using relabeling and relearning. Compared with Table II and III, we can know that the ACC further improves, and the ASR also decreases a little. Specifically, after employing relabeling and relearning, the average ACC and ASR changes from $84.29\%$ to $86.65\%$ and from $0.33\%$ to $0.03\%$ against the $8$ backdoor attacks, respectively. We infer that the model learns better due to the increase of clean samples. More interestingly, we find that employing relabeling and relearning has a positive effect on clean-label backdoor attacks (e.g., SIG). The ASR drops from $2.64\%$ to $0.22\%$ after equipping the relabeling and relearning mechanism. We deduce that the efficacy of backdoor task learning is closely linked to the initial weights of the model. Therefore, a pretrained model can hardly be backdoored by poisoned samples, which also indicates model service providers use pretrained models as the base models and fine-tune them with untrusted data to avoid poisoning attacks.
@@
Stratagies→ W/O Relabeling and Relearning W/ Relabeling and Relearning
@@
VI-C2 Effectiveness with Different Partition Rates
In our AdvrBD method, we have to select a certain percentage of samples based on the $\mathcal{L}_{kl}$ value for training a backdoor enhancement model. This means that the partition rate may be important for AdvrBD’s performance. Therefore, to explore the effect of partition rate on the performance, we conduct experiments on CIFAR10 and $8$ backdoor attacks when the partition rate is $0.05$, $0.10$, $0.15$, $0.20$, and $0.25$, respectively. The experimental results are shown in Fig.7, which show that the partition rates have little influence on the AdvrBD’s performance. Our method on all the backdoor attacks apart from SIG has a stable performance. The ASR is always nearly $0.00\%$ except for SIG, in which the ASR causes a little fluctuation with a very small margin around $0.00\%$. We infer that it’s a normal fluctuation because of the occasionality of each experiment. For ACC, there is also a little fluctuation around $85.00\%$ for all the backdoor attacks. We infer that it is also a normal fluctuation due to the occasionality of each experiment.
Figure 7: The ASR and ACC (%) of AdvrBD with different partition rates.
@@ VI-C2 Effectiveness with Different Partition Rates
Poison Ratio→ 0 0.1 0.15 0.2
Method↓ ACC ASR ACC ASR ACC ASR ACC ASR
BadNet 86.04 0.66 85.82 0.00 85.38 0.00 82.20 0.00
@@
Average 86.04 0.66 83.50 0.00 83.87 0.00 81.98 0.00
VI-C3 Effectiveness with Different Poisoning Ratios
In real-world application scenarios, it is impractical for the defender to access the poisoning ratios of training data. Therefore, we also demonstrate that our AdvrBD is suitable for multiple poison ratios. Here, we experiment it on CIFAR10 against $7$ backdoor attacks, including BadNet, Blend, WaNet, PhysicalBA, TaCT, AdaptivaBlend, and AdaptivaPatch with poisoning ratios up to $0.1$, $0.15$, $0.2$. Note that SIG is not considered here because it is a clean-label attack that can just poison one-class samples. Moreover, the number of one-class samples is a maximum of $10.00\%$ in CIFAR10. That is, the poisoning ratios of SIG on CIFAR10 do not exceed $10.00\%$. Table V shows the experimental results. From Table V, we can know that with the variation of poisoning ratios, the ASR of our method is always $0.00\%$ against the $7$ backdoor attacks. In particular, we can find that the ASR is $0.66\%$ when the poisoning ratio is zero. We suppose that the reason is the misclassification of the model on the clean samples. In addition, the experimental results show that our method also maintains a fine ACC against different backdoor attacks even without employing relabeling and relearning methods.
VI-C4 Effectiveness with Different Models
To demonstrate the generalizability of our method, we conduct experiments on ResNet34 and MobileNetV2 with the poisoning ratio of $0.05$. Table VI shows the experimental results on CIFAR10. From Table VI, we can find that our AdvrBD works similarly well on ResNet34 and MobileNetV2. The ASR is always $0.00\%$ in the two model architectures. Moreover, the ACC also maintains a satisfactory performance (from $88.60\%$ to $88.70\%$ in ResNet34, from $84.33\%$ to $81.89\%$ in MobileNetV2). The experimental results indicate that our AdvrBD has good generalizability across different models.
@@ VI-C4 Effectiveness with Different Models
VI-C5 Impact on clean and poisoned dataset
In real-world applications, it is impractical for the defender to know whether a given training dataset contains some poisoned samples. Therefore, to demonstrate the impact of our AdvrBD on the clean and poisoned dataset, we conduct experiments on CIFAR10 with the poisoning ratio of 00 and $0.05$. The experimental results (Table VII) illustrate that for training on the poisoned dataset, our AdvrBD can significantly reduce the average ASR (from $74.45\%$ to $0.33\%$) with a slight loss of ACC (from $85.81\%$ to $84.29\%$). On the other hand, when trained on a clean dataset, our AdvrBD has a minor impact on ACC (from $87.86\%$ to $86.04\%$). Note that the ASR is $0.66\%$ when training on the clean dataset with our AdvrBD. We infer that this can be attributed to the misclassification of the model on the clean samples. The experimental results further demonstrate that AdvrBD is a practical solution for dealing with unknown datasets.
TABLE VII: Impact of AdvrBD on the clean and poisoned dataset.
Attack Method Poison w/o defense Poison w/ defense Clean w/o defense Clean w/ defense
ACC ASR ACC ASR ACC ASR ACC ASR
BadNet 92.00 95.20 85.72 0.00 87.86 0.00 86.04 0.66
@@ VI-C5 Impact on clean and poisoned dataset
Average 85.81 74.45 84.29 0.33 87.86 0.00 86.04 0.66
VII Conclusion
In this paper, we first explore the relationship between the backdooor and perturbation by our theoretical analysis and experimental verification. Based on the exploration results, we propose a novel backdoor defense method (AdvrBD) to identify poisoned samples and train a clean model on a poisoned dataset. The proposed AdvrBD partitions the poisoned samples and clean samples based on their robustness to the adversarial perturbation. There is no requirement for any auxiliary clean dataset or knowledge about the poisoned dataset (e.g., poisoning ratios) in the AdvrBD. Extensive experimental results show the superior performance of AdvrBD in defending against $8$ state-of-the-art backdoor attacks. Compared with the $7$ advanced backdoor defense methods, our AdvrBD has a lower ASR while maintaining a satisfactory ACC on the main task. Besides, the experimental results also demonstrate that the proposed AdvrBD has fine generalization capability in different poisoning ratios and various model architectures. The relevant experimental results indicate our method has prominent potential and vital practicality for real-world application scenarios as well.

### v2

@@ 5.3 The Resistance to Adaptive Attacks
To further reveal the potential risk of the MeCa, we also consider the adaptive adversaries that try to design special backdoor attacks to escape our MeCa method. The adaptive adversaries deliberately establish dependencies between the backdoor and normal functionality, which immensely increases the difficulty of backdoor sample detection. Therefore, we employ three adaptive attacks (TaCT [43], AdaptiveBlend [49] and AdaptivePatch [49]) to demonstrate the effectiveness of our defense method. The three adaptive attacks build the dependencies between the backdoor and normal functions by using different poisoning strategies. Table III shows the defense results of our method and other $7$ state-of-the-art backdoor defense techniques against the $3$ adaptive attacks on $4$ different datasets. From Table III, we have the following three observations. First, our MeCa has the lowest ASR (nearly $0.00\%$) while maintaining a satisfactory ACC of the main task. Then, some existing defense methods (CLP and SCAnFT) have a good defense performance against the three adaptive attacks, but they have an obvious decrease in ACC. Specifically, for CLP on the CIFAR10, the ASR is $3.55\%$, and the ACC drops from $83.63\%$ to $52.56\%$. Finally, our method has a better generalization on different datasets. For different datasets, our method always maintains a fine performance, while the performance of the other backdoor defense techniques has an obvious fluctuation. In summary, compared with the $7$ state-of-the-art backdoor defense techniques, our method has a better performance and greater generalization against adaptive attacks.
5.4 Ablation Studies
@@
Different from the existing backdoor defense methods, we also propose to improve the performance of our MeCa further by relabeling and relearning the backdoor samples. Hence, in this section, we first show the MeCa’s performance after employing relabeling and relearning mechanisms. Then, to demonstrate our MeCa’s generalization, we also experimentally explore the performance of our MeCa on different partition rates, poison ratios, and models. Finally, we also explore the impact of our MeCa on the clean and poisoned dataset.
5.4.1 Performance with Relabeling and Relearning
@@
In our MeCa, after identifying the backdoor samples and clean samples, we relabel the backdoor samples and merge them with clean samples to obtain a clean and complete dataset, which is used to fine-tune the clean model then. Accordingly, we conduct experiments on the ResNet18 and CIFAR10 datasets to show the effect of the relabeling and relearning mechanism. Table IV shows the ACC and ASR of our MeCa after using relabeling and relearning. Compared with Table II and III, we can know that the ACC further improves, and the ASR also decreases a little. Specifically, after employing relabeling and relearning, the average ACC and ASR changes from $84.29\%$ to $86.65\%$ and from $0.33\%$ to $0.03\%$ against the $8$ backdoor attacks, respectively. We infer that the model learns better due to the increase of clean samples. More interestingly, we find that employing relabeling and relearning has a positive effect on clean-label backdoor attacks (e.g., SIG). The ASR drops from $2.64\%$ to $0.22\%$ after equipping the relabeling and relearning mechanism. We deduce that the efficacy of backdoor task learning is closely linked to the initial weights of the model. Therefore, a pretrained model can hardly be backdoored by poisoned samples, which also indicates model service providers use pretrained models as the base models and fine-tune them with untrusted data to avoid poisoning attacks.
@@
Stratagies→
W/O Relabeling
and Relearning
W/ Relabeling
and Relearning
@@
5.4.2 Effectiveness with Different Partition Rates
In our MeCa method, we have to select a certain percentage of samples based on the $\mathcal{L}_{kl}$ value to train a backdoor enhancement model. This means that the partition rate may be important for MeCa’s performance. Therefore, to explore the effect of partition rate on performance, we conduct experiments on CIFAR10 and $8$ backdoor attacks when the partition rate is $0.05$, $0.10$, $0.15$, $0.20$, and $0.25$, respectively. The experimental results are shown in Fig.7, which shows that the partition rates have little influence on the MeCa’s performance. Our method on all the backdoor attacks apart from SIG has a stable performance. The ASR is always nearly $0.00\%$ except for SIG, in which the ASR causes a little fluctuation with a very small margin around $0.00\%$. We infer that it’s a normal fluctuation because of the occasionality of each experiment. For ACC, there is also a little fluctuation around $85.00\%$ for all the backdoor attacks. We infer that it is also a normal fluctuation due to the occasionality of each experiment.
Figure 7: The ASR and ACC (%) of MeCa with different partition rates.
@@ 5.4.2 Effectiveness with Different Partition Rates
Poison Ratio→ 0.1 0.15 0.2
Method↓ ACC ASR ACC ASR ACC ASR
BadNet 85.82 0.00 85.38 0.00 82.20 0.00
@@
Average 83.50 0.00 83.87 0.00 81.98 0.00
  • $\bullet$ Without attack: ACC-86.04%, ASR-0.66%
5.4.3 Effectiveness with Different Poisoning Ratios
In real-world application scenarios, it is impractical for the defender to access the poisoning ratios of the training data. Therefore, we also demonstrate that our MeCa is suitable for multiple poison ratios. Here, we experiment it on CIFAR10 against $7$ backdoor attacks, including BadNet, Blend, WaNet, PhysicalBA, TaCT, AdaptivaBlend, and AdaptivaPatch with poisoning ratios up to $0.1$, $0.15$, $0.2$. Note that SIG is not considered here because it is a clean-label attack that can just poison one-class samples. Moreover, the number of one-class samples is a maximum of $10.00\%$ in CIFAR10. That is, the poisoning ratios of SIG on CIFAR10 do not exceed $10.00\%$. Table V shows the experimental results. From Table V, we can know that with the variation of poisoning ratios, the ASR of our method is always $0.00\%$ against the $7$ backdoor attacks. In particular, we can find that the ASR is $0.66\%$ when the poisoning ratio is zero. We suppose that the reason is the misclassification of the model on the clean samples. In addition, the experimental results show that our method also maintains a fine ACC against different backdoor attacks even without employing relabeling and relearning methods.
5.4.4 Effectiveness with Different Models
To demonstrate the generalizability of our method, we conduct experiments on ResNet34 and MobileNetV2 with the poisoning ratio of $0.05$. Table VI shows the experimental results on CIFAR10. From Table VI, we can find that our MeCa works similarly well on ResNet34 and MobileNetV2. The ASR is always $0.00\%$ in the two model architectures. Moreover, the ACC also maintains a satisfactory performance (from $88.60\%$ to $88.70\%$ in ResNet34, from $84.33\%$ to $81.89\%$ in MobileNetV2). The experimental results indicate that our MeCa has good generalizability between different models.
@@ 5.4.4 Effectiveness with Different Models
TABLE VII: Impact of MeCa on the poisoned dataset.
Poison w/o defense Poison w/ defense
Attack Method ACC ASR ACC ASR
BadNet 92.00 95.20 85.72 0.00
@@
Average 85.81 74.45 84.29 0.33
  • $\bullet$ Clean dataset without defense: ACC-87.86%, ASR-0.00%
  • $\bullet$ Clean dataset with defense: ACC-86.04%, ASR-0.66%
5.4.5 Impact on clean and poisoned dataset
In real-world applications, it is impractical for the defender to know whether a given training dataset contains some poisoned samples. Therefore, to demonstrate the impact of our MeCa on the clean and poisoned dataset, we conduct experiments on CIFAR10 with the poisoning ratio of 00 and $0.05$. The experimental results (Table VII) illustrate that for training on the poisoned dataset, our MeCa can significantly reduce the average ASR (from $74.45\%$ to $0.33\%$) with a slight loss of ACC (from $85.81\%$ to $84.29\%$). On the other hand, when trained on a clean dataset, our MeCa has a minor impact on ACC (from $87.86\%$ to $86.04\%$). Note that the ASR is $0.66\%$ when training on the clean dataset with our MeCa. We infer that this can be attributed to the misclassification of the model on the clean samples. The experimental results further demonstrate that MeCa is a practical solution for dealing with unknown datasets.
6 Conclusion
In this paper, we first explore the relationship between the backdooor and perturbation by our theoretical analysis and experimental verification. Based on the results of the investigation, we propose a novel backdoor defense method (MeCa) to identify poisoned samples and train a clean model on a poisoned dataset. The proposed MeCa partitions the poisoned samples and clean samples according to their robustness to adversarial perturbation. There is no requirement for any auxiliary clean dataset or knowledge about the poisoned dataset (e.g., poisoning ratios) in the MeCa. Extensive experimental results show the superior performance of MeCa in defending against $8$ state-of-the-art backdoor attacks. Compared with the $7$ advanced backdoor defense methods, our MeCa has a lower ASR while maintaining a satisfactory ACC on the main task. In addition, the experimental results also demonstrate that the proposed MeCa has a fine generalization ability in different poisoning ratios and various model architectures. The relevant experimental results indicate our method has prominent potential and vital practicality for real-world application scenarios as well.